1)    Which of the following are in O (n):

a) n + lg n        b) n + 2n        c) n + n2            d) 1000 n + 4500 lg n + 54 n

2)    What is a stack? Give an ADT.

3)    How would you implement the push method for a Stack implemented as a Linked List?

4)    How would you implement the pop method for a Stack implemented as a Linked List?
Answer #1
1)
a) n + lg n = O(n)
b) n + 2n = O(2^n)
c) n + n2 = O(n^2)
d) 1000 n + 4500 lg n + 54 n = O(n)

so, a and d are in O(n)